# WakenLLM: A Fine-Grained Benchmark for Evaluating LLM Reasoning Potential and Reasoning Process Stability

### v1

@@ Abstract
Large Language Models (LLMs) frequently output the label Unknown, yet current evaluations focus almost exclusively on whether such answers are honest rather than why they arise. This blurs two distinct cases: (i) an input that is genuinely indeterminate and (ii) a solvable problem that the model fails to resolve. We call this phenomenon Vague Perception. And thus we introduce a framework that quantifies the proportion of Unknown responses attributable to model incapacity and tests whether guided stimulation can convert them into either correct (Known) or intrinsically indeterminate outcomes. By separating these sources of uncertainty, our method provides a clearer picture of LLM reasoning limits and their potential for improvement. As we get a theoretical accuracy of reasoning task on different LLMs, we apply different methods to test whether the model can reach the accuracy given a baseline framework. Our work is meaningful in exploring the true reasoning ability of LLMs and providing a new perspective on solving the Vague Perception phenomenon.
@@ Introduction
Fine-grained reasoning reliability remains under-explored. Empirical studies have shown that LLMs frequently hallucinate, hedge, or output “Unknown” even when the context contains sufficient evidence, while at other times they confidently answer questions that are fundamentally under-specified (Lin 2023; Ye and et al. 2023). This paper formalizes this phenomenon as Vague Perception: a mismatch between a model’s internal reasoning process and the logical sufficiency of the input. Therefore, a critical gap lies in understanding how LLMs parse complex information, decide that evidence is (in)sufficient, and transition from partial inference to a definitive judgment when faced with ambiguity or uncertainty.
Figure 1: Definition of Vague Perception: when an LLM misjudges a Verifiable sample as (“Unknown”), or it is unable to tell why a sample should be categorized as unknown inputting Unverifiable Samples. The figure presents two representative examples attributed to distinct causes: Fact Understanding and Reasoning Gap—highlighting 1. Limited ability to solve a problem. 2.Misunderstandings of facts versus failures to link them.
In this work, we analyze the root causes of deep-reasoning failures behind Vague Perception and propose fine-grained mitigation strategies. To enable systematic measurement, we introduce a new benchmark that spans four subject domains and six subcategories, offered in two complementary formats: (i) Fact-based benchmark, where relevant information is listed as discrete atomic statements (e.g., “A is X, B is Y”); and (ii) Story-based benchmark, where the same facts are embedded in a coherent passage. This design lets us probe how presentation style and information density influence the emergence and eventual resolution of Vague Perception.
Paradigm Guidance Questioning Stimulation Dynamic Samples Selection Reflection Stage-wise Examining Fine-Grained Decomposing
@@
We construct benchmarks in which half of the samples has T/F labels and the other half is Unknown. These samples are then fed into the model to extract the corresponding Vague Perception samples. Subsequently, the selected instances were processed via a two-stage stimulation mechanism aimed at “waking up” the model’s latent reasoning abilities: In stage 1, we enumerate explicit rules that the LLM should follow in order to reanalyze the problems.
@@
We propose WakenLLM as a strong baseline to explore the reasoning abilities of LLMs among Vague Perception samples, yielding comparable performance with GPT-3.5. Furthermore, we evaluate the performance of over 30 LLMs across these reasoning methods on TableBench, highlighting that both open-source and proprietary LLMs require substantial improvements to meet real-world demands. Notably, even one of the most advanced models, GPT-4o, achieves only a modest score.
@@
  • We perform a comprehensive investigation of Vague Perception on both Fact-based and Story-based benchmarks, analyzing its influence on reasoning in casual and professional benchmarks.
  • We introduce WakenLLM, a unified framework to evaluate the ”Awaken Ability” of LLMs—their capacity to produce correct answers after stimulation—representing a principled approach to handling Vague Perception samples.
  • Through a systematic analysis of models with varying parameter scales, we establish benchmark metrics that more precisely quantify the challenges posed by Vague Perception. The resulting insights not only improve reasoning accuracy but also advance our understanding of LLM trustworthiness and in-depth reasoning paradigm.
@@
The WakenLLM evaluation framework, illustrated in Figure 2, is composed of two successive stages. It evaluates the model’s Awakening Ability and systematically addresses the challenge of Vague Perception. Specifically, the first stage assesses the Overall Benchmark Converting Rate, while the second investigates the impact of stimulation under the Remind-then-Guide (RtG) paradigm on both label and reasoning process conformity.
@@
Our framework begins with a model-specific pre-processing stage that identifies Vague Perception samples—cases where the LLM initially mislabels (e.g., saying Unknown when the sample is verifiable, or failing to explain an unverifiable case). Since this process is performed separately for each model, the set of Vague Perception samples differs across models. The converting results fall into three categories: (1) The model fails to generate a coherent logic chain, which we call Unexcited Converting (UC). (2) The model produces a false logic chain and an incorrect final answer, denoted as False Converting (FC). (3) The model outputs a correct label under stimulation, called True Converting (TC). To quantify these proportions, we use UCR, FCR, and TCR to represent the rates of each type of conversion. Given our focus on the genuine reasoning correction of the model, we elaborate only on TCR. UCR and FCR follow analogous logic. In order to measure the ability of judging both verifiable and unverifiable data, we denote 2 separate metrics for each, defining True Converting Rate(TCR) in N-th stage as follows:
$\begin{gathered}TC_{f}^{n}\;=\;\sum_{y_{f}\in D_{\mathrm{VP}}^{n}}\mathbf{1}\bigl{(}\hat{y}_{f}^{n}=y_{f}\bigr{)},\\[6.0pt] TCR_{f}^{n}\;=\;\frac{\bigl{|}TC_{f}^{n}\bigr{|}}{\bigl{|}D_{\mathrm{VP}}^{n}\bigr{|}}\,,\\[6.0pt] \text{where}\quad f\in\{u,v\},\quad n\in\{1,2\}.\end{gathered}$
Figure 2: Overview of the proposed pipeline. We begin with a balanced dataset with Fact and Story-based dataset, sampling verifiable (T/F) and unverifiable (Unknown) instances at a 1 : 1 ratio. The pipeline then (1) detects cases that expose the model’s Vague Perception during reasoning and (2) applies a two-stage remedy: Stage 1 uses stimulation to re-evaluate the mislabeled samples, while Stage 2 gathers the remaining False Converting samples from Stage 1 for model self-reflection. Stimulation Types design vary with data form.
Here, $TC_{v}$ quantifies input samples that have been correctly classified, while $TC_{u}$ refers to unverifiable samples where an Unknown label has been correctly assigned. $N$ denotes the stimulation stage, which can be either one or two. $D_{\text{VP}}^{(n)}$ represents the input of Vague Perception samples in the $n$-th stage. We can tell that $D_{\mathrm{VP}}^{2}=FC^{2}$.
It should be noted that both $TCR_{v}$ and $TCR_{u}$ quantities represent the results at the current stage, rather than the cumulative sum across both stages. To integrates performance across both stages, we define Overall Converting Rate (OCR) as follows:
@@
It refers to the portion of under-stimulated samples turn out to be predicted right. By doing this, we can obtain the theoretical potential accuracy output by a model.
@@
Label Offering
We test whether Vague Perception samples display a better or worse stability—i.e. whether they are more easily misled—relative to ordinary samples. According to previous work (Weng, Chen, and Wang 2025), all models demonstrate Conformity. In this part, we first inform the model that it has misunderstood the sample, and then provide a randomly assigned label with varying probabilities, to test whether this misdirection affects its final decision.
@@
X can be v or u, indicating whether a sample is objectively verifiable or unverifiable on its own. Here, $S$ denotes the model is under the RTG stimulation.
Reasoning Process (RP) Offering
We examine how effectively the reasoning chain generated in Stage 1 serves as inspiration or guidance for Stage 2 when it is supplied as additional input. To ensure that the model demonstrates genuine understanding rather than simply being prompted to output the correct answer by accident, we propose two metrics.
Correct Guiding Rate (CGR): This metric assesses whether the reasoning process output in Stage 1 can assist Stage 2 in stimulating samples to better avoid prior mistakes.
Reasoning Process Coherence (RPC): This metric involves adding the reasoning process from Stage 1 to re-run Stage 1, measuring the alignment of the reasoning process provided by Stage 1 to assess its stability. where R1 denotes the input contains the reasoning process output by Stage 1.
@@
To explore the effects of the Vague Perception phenomenon on different types of logical data, we classify them based on their labels: (1) Verifiable Data, where the samples have a clear True/False label indicating the validity of the reasoning; and (2) Unverifiable Data, where the label is Unknown and samples are inherently unsolvable.
Furthermore, as shown in Figure 3, to better understand how large language models (LLMs) reason across different logical forms, we categorize the whole benchmarks as follows:
@@
FLD (Morishita et al. 2024) supplies three kinds of sample labels annotated as True, False, or Unknown. We randomly sample 600 items: 300 verifiable (True/False) and 300 marked Unknown due to insufficient evidence. FOLIO (Han et al. 2022) is a factual benchmark with the same 3-way labeling. Following an identical protocol, we draw 640 instances—320 verifiable claims and 320 Unknown.
Figure 3: Overall benchmark distribution. Our dataset contains 2,840 samples grouped into two top-level task types: story-based (56.3%) and fact-based (43.7%). A detailed breakdown of the “Else” category is provided in the Appendix.
@@
For ScienceQA (Lu et al. 2022) we construct two balanced splits to capture both Science and Arts reasoning. The Science split contains 400 Physics and 400 Biology questions; the Arts split holds 400 Figurative-Language and 400 Writing-Strategies questions. Within each split exactly half of the items are answerable from the passage, while the remaining half are deliberately labelled Unknown, mirroring the fact-verification setting and enabling a unified evaluation protocol.
Our experimental design considers the differing characteristics of Fact-based benchmark and Story-based benchmark. For fact-based tasks, which models tend to process the entire context holistically, so providing additional stimulation can enhance performance. In contrast, for story-based tasks—often rooted in professional knowledge domains and relying on models’ inherent knowledge—we minimize guidance to avoid interfering with the model’s independent reasoning process.
Experimental Setup
Our evaluation covers both closed-source (ChatGPT-3.5-turbo and GPT-4o) and open-source models (Llama-3.1, Gemma-2, and Qwen-2.5 series) (OpenAI 2022; Touvron and Team 2025; Google DeepMind 2025; Bai and Team 2025). Two stimulating strategies are implemented:
  • Detailed Stimulation: Enhanced detailed approach featuring:
    • Listed Conditions Guidance
    • Structured Reasoning Constraints
    • Conditional Verification Steps
The experimental protocol differs between approaches. Concise Stimulation uses a similar format for two-stage error notification process, while Detailed Stimulation employs a different two-stage workflow: a standard prediction followed by an interrogative refinement. The latter stimulates analytical reasoning through rhetorical questioning and step-by-step error analysis, creating a clear methodological contrast.
Results on Overall Benchmark Converting Rate
As shown in Table 3, we report the overall benchmark Converting Rate, which means 4 metrics in three systems: Unexcited Converting Rate (UCR), False Converting Rate (FCR), and True Converting Rate for both verifiable and unverifiable data (TCRv, TCRu).
Benchmark Samples Target Form Writing Type
@@
Table 3: Comparison of six LLMs across four task subclasses (FLD, FOLIO, Science, Arts), grouped by evaluation settings.
@@
Label Conformity
Judging from the table, we observe that even when LLMs are reminded of their previous incorrect answers, they still demonstrate a strong tendency toward conformity. Moreover, distinct conformity patterns emerge between verifiable and unverifiable samples. One key influencing factor appears to be the model size: models with smaller parameter sizes, such as GPT-3.5 and LLaMA-3.1-8B, are more prone to being misled on verifiable samples involving T/F label reasoning. In contrast, models with larger parameter sizes, such as GPT-4o and LLaMA-3.1-405B, tend to be more easily misled when the label is Unknown. Meanwhile, Qwen series shows a distinctive feature different from the above two kinds, we infer it is because parameters gap between Qwen is much smaller than those.
Reasoning Process Conformity
When we input the previous reasoning process and explicitly ask the LLM to reflect on its mistakes, the resulting CGR values are mostly below zero. This indicates that even after being reminded of a past error and provided with its own flawed reasoning, the LLM still tends to repeat the same incorrect conclusion, leading to a noticeable decline in accuracy.
Furthermore, we investigate whether the reasoning process generated by LLMs is generally stable. Our results reveal that the reasoning process is often highly unstable. Changes in accuracy can largely be attributed to differences in how the problem is interpreted. Notably, for unverifiable samples, the reasoning process tends to be unreliable — feeding it back into the model typically leads to a further drop in accuracy. This exposes a hidden issue: when LLMs automatically generate solutions using a reasoning process, unverifiable problems are particularly susceptible to unreliability. Even for verifiable samples, the reasoning may remain unstable.
Discussion
As shown in the line chart of Fig 4, by comparing Stage 1 and 2 True Converting Rate (TCR1 and TCR2) across different benchmark types, we observe clear divergences: Fact-based benchmarks tend to achieve a markedly higher TCR2, implying that a great amount of mis-predicted samples got predicted right in Stage 2, where the model can effectively refine the outputs produced in Stage 1. In contrast, Story-based benchmarks’ converting rate excel in Stage 1 but show a substantially lower TCR2. Thus, distinct stimulation strategies are recommended: For fact-based benchmarks, broader exploratory freedom is acceptable in Stage 1. “Making mistakes at first is allowed”, since later refinement proves beneficial; For Story-based benchmarks, the model must “do it right the first time,” as the improvement of subsequent stage is limited. Hence, optimization priorities diverge: Fact-based tasks profit from a two-stage pipeline that revises early missteps, whereas story-based tasks demand maximizing single-pass accuracy.
Detailed Analysis
Benchmark-wise Vague Perception

### v2

# WakenLLM: Evaluating Reasoning Potential and Stability in LLMs via Fine-Grained Benchmarking

@@ Abstract
Large Language Models (LLMs) frequently output the label Unknown, yet current evaluations focus almost exclusively on whether such answers are honest rather than why they arise. This blurs two distinct cases: (i) an input that is genuinely indeterminate and (ii) a solvable problem that the model fails to resolve. We call this phenomenon Vague Perception. And thus we introduce a framework that quantifies the proportion of Unknown responses attributable to model incapacity and tests whether guided stimulation can convert them into either correct (Known) or correct Unknown with valid reasoning. By separating these sources of uncertainty, our method provides a clearer picture of LLM reasoning limits and their potential for improvement. As we get a theoretical accuracy of reasoning task on different LLMs, we apply different methods to test whether the model can reach the accuracy given a baseline framework. Our work is meaningful in exploring the potential reasoning ability of LLMs and providing a new perspective on solving the Vague Perception phenomenon.
@@ Introduction
Fine-grained reasoning reliability remains under-explored. Empirical studies have shown that LLMs frequently hallucinate, hedge, or output “Unknown” even when the context contains sufficient evidence, while at other times they output Unknown without understanding why, facing fundamentally under-specified (Lin 2023; Ye and et al. 2023), (Zhou et al. 2025) problems. This paper formalizes this phenomenon as Vague Perception: a mismatch between a model’s reasoning ability and the objective problems. Therefore, a critical gap lies in understanding how LLMs parse complex information, decide that evidence is (in)sufficient, and transition from uncertainty inference to a definitive judgment when faced with hard unsolvable or uncertain problems.
Figure 1: Definition of Vague Perception: when an LLM misjudges a Verifiable sample as Unknown, or it is unable to tell why a sample should be categorized as unknown inputting Unverifiable Samples. The figure presents two representative reasons attributed to this phenomenon: Fact Understanding and Reasoning Gap—highlighting 1. Limited ability to solve a problem. 2.Misunderstandings of facts versus failures to link them.
In this work, we analyze the root causes of deep-reasoning failures behind Vague Perception and propose fine-grained mitigation strategies. To enable systematic measurement, we introduce a new benchmark that spans four subject domains and multiple subcategories, offered in two different formats: (i) Fact-based benchmark, where relevant information is listed as discrete atomic statements (e.g., “A is X, B is Y”); and (ii) Story-based benchmark, where the same facts are embedded in a coherent passage. This design lets us probe how presentation style and information density influence the emergence and eventual resolution of Vague Perception.
Paradigm Guidance Questioning Stimulation Dynamic Samples Selection Reflection Stage-wise Examining Fine-Grained Decomposing Diversity
@@
Trace-of-Thought (McDonald and Emami 2025)
MAPS (de Souza Loureiro et al. 2025)
CoD (Xu et al. 2025)
ReSearch (Chen et al. 2025)
@@
We construct benchmarks in which half of the samples are verifiable: with T/F labels and the other half is unverifiable: with Unknown labels that are objectively unsolvable. These samples are then fed into the model to extract the corresponding Vague Perception samples. As shown in Figure 1, whenever model output unknown without a clear understanding of the sample, we consider it to be vague perception. Subsequently, the selected samples were processed via a two-stage stimulation mechanism aimed at “waking up” the model’s reasoning abilities: In stage 1, we enumerate explicit rules that the LLM should follow in order to reanalyze the problems.
@@
We propose WakenLLM as a strong baseline to explore the reasoning abilities of LLMs among Vague Perception samples, yielding comparable performance with over five various LLMs across these reasoning tasks, highlighting that both open-source and closed-source LLMs show substantial improvements on reasoning accuracy.
@@
  • We discover Vague Perception, a phenomenon in reasoning tasks To examine it comprehensively, we propose WakenLLM, a unified framework that quantifies an LLM’s Awaken Ability—its capacity to convert to the correct answer once stimulated—offering a principled solution for handling Vague Perception samples in reasoning tasks.
  • Through systematic analyses across models of varying parameter scales, we devise benchmark metrics that more precisely characterize the challenges posed by Vague Perception. The resulting insights not only boost reasoning accuracy but also deepen our understanding of LLM trustworthiness and advanced reasoning paradigms.
  • By comparing WakenLLM against diverse baselines, we measure how effectively each approach unlocks an LLM’s latent potential. Incorporating both training and non-training frameworks, our work offers a new perspective on mitigating the Vague Perception phenomenon.
Figure 2: Overview of the proposed pipeline. We begin with a balanced dataset with Fact and Story-based dataset, sampling verifiable (T/F) and unverifiable (Unknown) samples at a 1 : 1 ratio. The pipeline then (1) detects cases that expose the model’s Vague Perception during reasoning and (2) applies a two-stage remedy: Stage 1 uses stimulation to re-evaluate the mislabeled samples, while Stage 2 gathers the remaining False Converting samples from Stage 1 for model self-reflection. Stimulation Types design vary with data form.
@@
The WakenLLM evaluation framework, illustrated in Figure 2 part 2, is composed of two successive stages. It evaluates the model’s potential reasoning ability and systematically evaluate the challenge of Vague Perception. For the vanilla setting, the first stage assesses the , while the second investigates the impact of stimulation, thus we can acquire a Overall Benchmark Converting Rate, indicating the percentage of Vague Perception samples can be reasoned right by LLMs. Under the Remind-then-Guide (RtG) setting, aside from reminding it did wrong previously, we assign label or previous reasoning process as the input, we examine whether LLMs still show conformity after reminding it solving the problem wrong previously, or it would learn from it and distinguish whether it is guiding or mis-guiding.
@@
Our framework begins with a model-specific pre-processing stage that identifies Vague Perception samples—cases where the LLM initially mislabels (e.g., saying Unknown when the sample is verifiable, or failing to explain an unverifiable sample). Since this process is performed separately for each model, the set of Vague Perception samples differs across models.
Gaining Vague Perception samples as the input of pipeline stage 1, the converting results fall into three categories: (1) The model fails to generate a coherent logic chain and remains output Unknown, which we call Unexcited Converting (UC). (2) The model produces a false logic chain and an incorrect final label, denoted as False Converting (FC). (3) The model outputs a correct label under stimulation, called True Converting (TC). To quantify these proportions, we use UCR, FCR, and TCR to represent the rates of each type of conversion. Given our focus on the reasoning correction of the model, we elaborate only on TCR and OCR. defining True Converting Rate(TCR) in n-th stage as follows:
$TC_{f}^{n}\;=\;\sum_{y_{f}\in D_{\mathrm{VP}}^{n}}\mathbf{1}\bigl{(}\hat{y}_{f}^{n}=y_{f}\bigr{)},\quad TCR_{f}^{n}\;=\;\frac{\bigl{|}TC_{f}^{n}\bigr{|}}{\bigl{|}D_{\mathrm{VP}}^{n}\bigr{|}}\,$
In the above equations, $f\in\{u,v\}$ denotes the sample type (verifiable or unverifiable), and $n\in\{1,2\}$ indicates the stimulation stage. $D_{\text{VP}}^{(n)}$ represents the input of Vague Perception samples in the $n$-th stage. First stage input $D_{\mathrm{VP}}^{1}$ was extracted from raw dataset. For the second stage, we collect the False Converting part samples as the input, so we have: $D_{\mathrm{VP}}^{2}=FC^{2}$.
It should be noted that both $TCR_{v}$ and $TCR_{u}$ quantities represent the results sum both stage, and $TCR^{1}$ denotes the converting rate of both forms data in first stage. To be specify, $TCR_{v}$ = $TCR_{v}^{1}$ + $TCR_{v}^{2}$, and $TCR^{1}$ = $TCR_{v}^{1}$ + $TCR_{u}^{1}$. To integrates performance across both stages, we define Overall Converting Rate (OCR) as follows:
@@
OCR refers to the portion of under-stimulated samples turn out to be predicted right. By doing this, we can obtain the theoretical potential accuracy output by a model.
@@
Label Conformity
We test whether Vague Perception samples display a better or worse stability—i.e. whether they are more easily misled—relative to ordinary settings indicated in (Weng, Chen, and Wang 2025). According to previous work, all models demonstrate Conformity. In this part, we first inform the model that it has misunderstood the sample, and then provide a randomly assigned label with varying probabilities, to test whether this misdirection affects its final decision.
@@
X can be v or u, indicating whether a sample is objectively verifiable or unverifiable on its own. The final result is average of both stage results.
Reasoning Process (RP) Conformity
We examine how effectively the reasoning chain generated in Stage 1 serves as inspiration or guidance for Stage 2 when it is supplied as additional input. To ensure that the model demonstrates genuine understanding rather than simply being prompted to output the correct answer by accident, we propose two metrics to dive deeper into this.
Correct Guiding Rate (CGR): CGR assesses whether the reasoning process output in Stage 1 can assist Stage 2 in stimulating samples to better avoid prior mistakes. Similarly, we remind the LLMs these may contain flaws since the samples are from False Converting parts. CGR is calculated by the changes of $TCR^{2}$ with and without Reasoning Process. The formula can be presented as: $\mathrm{CGR}=\triangle{\mathrm{TCR}}^{2}$
Reasoning Process Coherence (RPC): RPC involves adding the reasoning process from Stage 1 to re-run Stage 1, measuring the alignment of the reasoning process provided by Stage 1 to assess its stability. Similary, RPC is calculated by changes of $TCR^{1}$ with and without Reasoning Process. The formula can be presented as: $\mathrm{RPC}=\triangle{\mathrm{TCR}}^{1}$.
For both metrics, RPC and CGR, below zero indicates that the previous reasoning process is misleading and results in declined accuracy, while above zero suggests that it is helpful and can guide the LLM towards correct reasoning.
@@
To explore the effects of the Vague Perception phenomenon on different types of logical data, we classify them based on their labels: (1) Verifiable Data, where the samples have a clear True/False/Unknown label indicating the validity of the reasoning; and (2) Unverifiable Data, where the label is Unknown and samples are inherently unsolvable.
Furthermore, as shown in Figure 3, to better understand how LLMs reason across different logical forms, we categorize the whole benchmarks as follows:
@@
The information is given in atomic relationships, split by unrelated sentences.: FLD (Morishita et al. 2024) supplies three kinds of sample labels annotated as True, False, or Unknown. We randomly sample 600 items: 300 verifiable (True/False) and 300 marked Unknown due to insufficient evidence. FOLIO (Han et al. 2022) is a factual benchmark with the same 3-way labeling. Following an identical protocol, we draw 640 samples—320 verifiable samples with T/F labels and 320 Unknown.
Figure 3: Overall benchmark distribution, which contains 2,840 samples grouped into two forms samples: story-based (56.3%) and fact-based (43.7%). A detailed breakdown of the “Else” category is provided in the Appendix.
@@
The information is given in a paragraph form, sentences in the context might be directly associtate with each other, closer to daily dialogue.
For ScienceQA (Lu et al. 2022) we construct two balanced splits to capture both Science and Arts reasoning. The Science split contains 400 Physics and 400 Biology questions; the Arts split holds 400 Language and 400 Writing questions. Within each split exactly half of the items are verifiable from the passage, while the remaining half are deliberately labeled Unknown, mirroring the Fact-based setting and enabling a unified evaluation protocol.
Benchmark Samples Form Writing Type
@@
Our experimental design considers the differing characteristics of the two forms above. For Fact-based ones, which models tend to process the entire context holistically, so providing additional stimulation can enhance performance. In contrast, for story-based tasks—often rooted in professional knowledge domains and relying on models’ inherent knowledge—we minimize guidance to avoid interfering with the model’s independent reasoning process.
Experimental Setup
Our evaluation covers both closed-source (ChatGPT-3.5-turbo and GPT-4o) and open-source models (Llama-3.1 series, and Qwen-2.5 series) (OpenAI 2022; Touvron and Team 2025; Bai and Team 2025). Two stimulating strategies are implemented:
  • Detailed Stimulation: Enhanced detailed approach featuring: (1)Listed Conditions Guidance. (2)Structured Reasoning Constraints. (3)Conditional Verification Steps.The experimental protocol differs between approaches. Concise Stimulation uses a similar format for two-stage error notification process, while Detailed Stimulation employs a different two-stage workflow: a standard prediction followed by an interrogative refinement. The latter stimulates analytical reasoning through rhetorical questioning and step-by-step error analysis, creating a clear methodological contrast.
@@
Table 3: Comparison of six LLMs across four task subclasses (FLD, FOLIO, Science, Arts), grouped by evaluation settings. RP refers to Reasoning Process under Remind-then-Guide setting.
Results on Overall Benchmark Converting Rate
As shown in Table 3, we report the overall benchmark Converting Rate, which means three metrics in three systems: True Converting Rate (TCR) for both stage, and Overall Converting Rate (OCR) taking the portion of Vague Perception samples into consideration. The LLaMA and Qwen series show superior Overall Converting Rate (OCR) on small-parameter series, whereas the GPT series shows the opposite trend. We infer that, in general, higher parameter counts correlate with lower OCR. Situation may change attributed to the distinct architecture of GPT-4o.
@@
Label Conformity
Judging from the table, we observe that even when LLMs are reminded of their previous incorrect answers, they still demonstrate a strong tendency toward conformity. Moreover, distinct conformity patterns emerge between verifiable and unverifiable samples. One key influencing factor appears to be the model size: models with smaller parameter sizes, such as GPT-3.5 and LLaMA-3.1-8B, are more prone to being misled on verifiable samples with T/F labels reasoning. In contrast, models with larger parameter sizes, such as GPT-4o and LLaMA-3.1-405B, tend to be more easily misled regarding unverifiable samples with Unknown label. Meanwhile, Qwen series shows a distinctive yet somehow similar features different from the above two kinds, we infer it is because parameters gap between Qwen is much smaller than the previous series.
Reasoning Process Conformity
When we input the previous reasoning process and explicitly ask the LLM to reflect on its mistakes, the resulting CGR values are mostly below zero. This indicates that even after being reminded of a past error and provided with its own flawed reasoning, the LLM still tends to repeat the same incorrect conclusion, leading to a even noticeable decline in accuracy.
Our results on RPC reveal that the reasoning process is often highly unstable. Changes in accuracy can largely be attributed to differences in how the problem is interpreted. Notably, for unverifiable samples, the reasoning process tends to be unreliable — feeding it back into the model typically leads to a further drop in accuracy. This exposes a hidden issue: when LLMs automatically generate solutions using a reasoning process, unverifiable problems are particularly susceptible to unreliability. Even for verifiable samples, the reasoning may remain unstable.
Discussion
As shown in the line chart of Figure 4, by comparing Stage 1 and 2 True Converting Rate (TCR1 and TCR2) across different benchmark types, we observe clear divergences: For GPT-4o, Fact-based benchmarks tend to achieve a markedly higher TCR2, implying that a great amount of mis-predicted samples got predicted right in Stage 2, where the model can effectively refine the outputs produced in Stage 1. In contrast, Story-based benchmarks’ converting rate excel in Stage 1 but show a substantially lower TCR2. Whereas LLaMA-3.1-8B demonstrates much higher True Converting Rate in stage 1.
Thus, distinct stimulation strategies are recommended for above models: (i) For GPT-4o, when facing Fact-based benchmarks, broader exploratory freedom is acceptable in Stage 1. “Making Mistakes at First is Allowed”, since later refinement proves beneficial; For Story-based benchmarks, the improvement of subsequent stage is limited, so initial accuracy should be underlined (ii) For LLaMA-3.1-8B, it is always better to “Do it Right the First Time” among all four dataset.
[…]
Detailed Analysis
Benchmark-wise Vague Perception